Determine the temperature (in ∘C) of the gases in the following containers.

Container B, with volume 0.20 m3 , holds 2.0 mol of argon at a gauge pressure of 0.50 atm .
Answer #1

We can apply Ideal gas equation to calculate the temperature.

P = Pgauge + Patm = 0.5 + 1 = 1.5 atm

V = 0.2 m3 = 0.2*1000 L = 200 L

Ideal Gas Equation, PV=nRT

T = PV/nR = (1.5*200)/(2*0.0821)

T = 1827 K

T = 1827 - 273 oC

T = 1554 oC

I hope i was able to answer your question. If you need anymore explanation, please ask in comments. I would be happy to help.

Please rate the solution if you are satisfied with the solution.